40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Jul 6 16:44:37 2011, steve, kiwamu, Update, SUS, BS oplev spectra BS.jpg
    Reply  Wed Jul 6 18:15:26 2011, kiwamu, Update, SUS, Re: BS oplev spectra 
    Reply  Tue Jul 12 17:30:24 2011, steve,, Update, SUS, BS oplev spectra BS_oplev.jpg
       Reply  Thu Jul 14 09:38:50 2011, steve,, Update, SUS, BS oplev spectra BS_oplev.jpg
          Reply  Thu Jul 14 15:27:08 2011, steve,, Update, SUS, SUS oplev spectras 7x
             Reply  Fri Jul 15 09:25:02 2011, rana, Update, SUS, SUS oplev spectras 
Message ID: 4966     Entry time: Thu Jul 14 09:38:50 2011     In reply to: 4963     Reply to this: 4967
Author: steve, 
Type: Update 
Category: SUS 
Subject: BS oplev spectra 

Quote:

Quote:

Healthy BS oplev

 I repeated the BS oplev spectrum today and I do not understand why it does look different. I did it as Kiwamu describes it in entry#4948  The oplev servo was left ON!

 It is working today! Finally I repeated the BS spectra, that we did with Kiwamu last week

Attachment 1: BS_oplev.jpg  103 kB  | Hide | Hide all
BS_oplev.jpg
ELOG V3.1.3-